OpenStudy (anonymous):

If we want the inequalities x>−1 and |−1|>|x| to both be true, what is a possible value for x?

OpenStudy (anonymous):

IxI <1

OpenStudy (anonymous):

actually, it is x=1/2, x=-1/2, and/or x=0. But thank's!

OpenStudy (anonymous):

x=.99999 no?

OpenStudy (anonymous):

Nope. Not exactly. But thanks!

OpenStudy (anonymous):

x>−1 and |−1|>|x| x=.9999 .999>-1 and I-1I> .9999